A 3 litre, four-stroke, four-cylinder CI engine running at 3000 rpm produces 48 kW of brake power. Brake thermal efficiency is 30% and the calorific value of fuel is 42 MJ/kg. Assume density of air at ambient condition is 1.1.81 kg/m3, air-fuel ratio 21:1 and mechanical efficiency 0.8. Determine:

(a) Rate of fuel flow into the engine (kg/s)

(b) Rate of air flow into the engine (kg/s)  

(c) Mass rate of exhaust flow (kg/s)

(d) Volumetric efficiency

(e) Indicated thermal efficiency
Answer #1

solution - 0 na = - ... me = b. p b.pl maxcv . 48 _ = 3.81x16% kysee hxc.v 0.3 X 42000 ma = 21 -3 ing ma = 21 x = 21x 3.81 xni = iP X CV = me . 60 3.81x103 x 42000 ni = 0.375 = 3754 volumetric efficiency, in = i (m.) actual 1 (matotal Total swept vo